Write an equation of the line containing the given point and parallel to the given point express your answer in the form y=m+b
(-7,9) 5x=7y+6
The equation of the line is y = ?

Answer by nerdybill(7384)

Write an equation of the line containing the given point and parallel to the given point express your answer in the form y=m+b
(-7,9) 5x=7y+6
first, we need to find the slope of:
5x=7y+6
do this, by putting it in "slope-intercept" form:
5x=7y+6
5x-6=7y
(5/7)x - 6/7 = y
the slope of 5/7
so, a line that is parallel must have the same slope.
plug the slope (5/7) and the given point (-7,9) into the "point-slope" form:
y - y1 = m(x - x1)
y - 9 = (5/7)(x - (-7))
y - 9 = (5/7)(x + 7)
y - 9 = (5/7)x + 5
y = (5/7)x + 14
.
The equation of the line is y = (5/7)x + 14